OpenStudy (anonymous):

Snowy's Snow Cones has a special bubble gum snow cone on sale. The cone is a regular snow cone that has a spherical piece of bubble gum nested at the bottom of the cone. The radius of the snow cone is 2 inches, and the height of the cone is 3 inches. If the diameter of the bubble gum is 0.5 inches, which of the following can be used to calculate the volume of the cone that can be filled with flavored ice? 1 over 3(3.14)(2^2)(3) − 4 over 3(3.14)(0.253) 1 over 3(3.14)(3^2)(2) − 4 over 3(3.14)(0.253) 1 over 3(3.14)(2^2)(3) − 4 over 3(3.14)(0.53) 1 over 3(3.14)(3^2)(2) − 4 over 3(3.14)(0.53)

OpenStudy (anonymous):

please help

OpenStudy (danjs):

The volume of space left in the cone after the gum is placed, is the total cone volume take away the sphere volume

OpenStudy (anonymous):

lol i am heaving one of those days i think i figured it out

OpenStudy (anonymous):

i think its the third option

OpenStudy (danjs):

looking at the answers helps out too, they are all the same form, to give an idea

OpenStudy (anonymous):

thanks for the help though

OpenStudy (danjs):

Volume Cone - Volume Sphere = (1/3 * pi * r^2 * h ) - (4/3 * pi * R^3) r = cone radius h = cone height R=sphere radius , half the diameter

OpenStudy (danjs):

the sphere radius, 0.25 , does not match any answers, R^3 = 0.25^3 = 0.0156 unless you forgot the (3) cube at the end, and they are using 0.253 as the radius ,half 0.5

OpenStudy (anonymous):

0.25^3